nvestment A has a beta of 1.3 and an expected rate of return of 17.2%. Investment B has a beta of 0.8 and an expected rate of return of 11.9%. What is the equity premium (market risk premium)?
Answer #1

CAPM

r = Rf + Beta * Market Risk Premium

Investment A

0.172 = Rf + 1.3 * Market Risk Premium (Eq 1)

Investment B

0.119 = Rf + 0.8 * Market Risk Premium (Eq 2)

Subtract Eq 2 from Eq 1

0.172 - 0.119 = Rf + 1.3 * Market Risk Premium - Rf - 0.8 * Market Risk Premium

0.053 = 0.5 * Market Risk Premium

Market Risk Premium = 0.053/0.5

Market Risk Premium = 0.106

Market Risk Premium = 10.6%

Can you please upvote? Thank You :-)